# Performance of a SiPM-based, plastic scintillator muon veto prototype for CUPID

### v1

@@ 5 Discussion
Comparing the overall results of all prototypes, the swirly patterns of the WLS fiber used in the Small Panel and Prototypes 1 and 2 resulted in modules with high uniformity, light yield, and position reconstruction. Prototype 1 had the highest light yield and the capability to reconstruct the position of muon events to an area of 2525 cm2. Modules of 10050 cm2 would be simpler to install and operate than 2525 cm2 panels. Modules with multiple fibers and SiPMs in the same continuous volume provide some contingency against broken SiPMs as the modules can still operate well with one SiPM not functioning. Prototype 2 displayed a reduced gamma background by using coincidence triggering in a double panel module, but increased the number of channels by a factor of two. Taking out the position requirement will further reduce the number of channels per panel. While this design has an adequate detection efficiency, an improved factor could be desired for ultra-low background experiments. A way to achieve this is to increase the thickness of the panels, yielding a greater gamma and muon peak separation and therefore a better discrimination power. Another option is to use two stacked panels for coincidence triggers, which would considerably decrease the detection threshold as it is unlikely that one gamma event traverses both panels and thus increase the muon detection efficiency. Both options are currently being explored and optimized with the help of the information obtained from this study.
@@ 6 Conclusions
In this work, we designed, constructed, and characterized two novel prototypes for CUPID’s compact muon veto system. The proposed prototypes consisting of wavelength-shifting fiber-embedded plastic scintillators coupled to silicon photomultipliers demonstrated a high and uniform light yield and a muon tagging efficiency of 98%. These features will be critical to reduce muon-induced background and enable the CUPID experiment to reach a sensitivity level commensurate with its science goal. Additionally, these panels had efficient position reconstruction capabilities within an area of 2525 cm2. This is the first muon veto design that explores position reconstruction within the same panel using a swirly geometry for further decreasing backgrounds.
@@
These prototypes successfully provided high efficiency in muon detection and position reconstruction. Future efforts will focus on optimizing a coincidence-based approach to further suppress the false-positive rate and adapting the system for large-scale deployment in CUPID. Additional work still needs to be performed before the final CUPID muon veto design can be realized, including studies of prototype performance and longevity under the environmental conditions at the Laboratori Nazionali del Gran Sasso and consideration of any systematic effects that may arise from the muon distribution at Gran Sasso. In future work, we plan to address these uncertainties and the overall impact of all studies on the final CUPID muon veto design.
@@ Acknowledgments
We thank the technical staff of Wright Laboratory, particularly Tom Barker for his assistance in designing PCBs. We acknowledge the contributions of Penny Slocum, James Nikkel, Thomas Barker, Caitlin Gainey, Gabriel Hoshino, Jacky Hua, Neal Ma, Din-Ammar Tolj, Iffat Zarif, and Andrew Zheng to this research. The authors are appreciative of the members of the CUORE and CUPID collaborations for their support and expertise in this work. This paper includes data from a panel produced for a muon veto at the Low Background Counting Facility at the Lawrence Berkeley National Laboratory. We thank Rick Norman for providing this panel and feedback on this manuscript. We also thank Vincenzo Caracciolo for useful discussions and comments on the manuscript. Eljen Technologies, particularly Chris Maxwell, provided samples and extensive knowledge for this project, for which we are grateful. Kuraray and Saint-Gobain (now named Luxium Solutions) provided fiber samples and expertise. We would also like to thank Hamamatsu Photonics and CAEN Technologies for their technical support, particularly Yuri Venturini and Marco Locatelli. This work is supported by the DOE Office of Science, Office of Nuclear Physics, Contract No. DE-SC0012654. This work made use of pandas [43], Matplotlib [44], Numpy [45], SciPy [46], and Jupyter [47].

### v2

@@ 5 Discussion
Comparing the overall results of all prototypes, the swirly pattern of the WLS fiber used in the Small Panel and Prototypes 1 and 2 resulted in modules with high uniformity, light yield, and position reconstruction. Prototype 1 had the highest light yield and the capability to reconstruct the position of muon events to an area of 2525 cm2. Modules with a surface area of 10050 cm2 would be simpler to install and operate than 2525 cm2 panels. Additionally, modules with multiple SiPMs connected in parallel in the same continuous volume provide some contingency against broken SiPMs. Operating a Prototype 1 panel with only 7 SiPMs instead of 8 results in a detection efficiency of , constituting a 6% decrease from the fully operational panel (see section 4.4).
From the tested configuration (directly stacked 1 cm panels), it was not evident that Prototype 2 would significantly decrease the gamma background, which would result in an increase of the detection efficiency and a decrease of the false-positive rate. Adding separation between two thin panels (Prototype 1), either air or another material such as lead, could potentially achieve this result. However, the cost would be an increase in the size of each module, which could conflict with CUPID’s design constraints. Prototype 2 would also increase the number of channels by a factor of two. However, taking out the position requirement will further reduce the number of channels per panel.
While Prototype 1 has an adequate detection efficiency, an improved factor could be desired for ultra-low background experiments. A way to achieve this is to increase the thickness of the panels, yielding a greater gamma and muon peak separation and therefore a better discrimination power. These options for Prototypes 1 and 2 are currently being explored and optimized with the help of the information obtained from this study.
@@ 6 Conclusions
In this work, we designed, constructed, and characterized two novel prototypes for CUPID’s compact muon veto system. The proposed prototypes consisting of wavelength-shifting fiber-embedded plastic scintillators coupled to silicon photomultipliers demonstrated a high and uniform light yield and a muon tagging efficiency of 98%. These features will be critical to reduce muon-induced background and enable the CUPID experiment to reach a sensitivity level commensurate with its science goal. Additionally, these panels had efficient position reconstruction capabilities within an area of 2525 cm2. This is the first muon veto design that explores position reconstruction within a single panel using a swirly geometry. The position reconstruction capability could be useful for further decreasing backgrounds.
@@
These prototypes successfully provided high efficiency in muon detection and position reconstruction. Future efforts will focus on optimizing a coincidence-based approach to further suppress the false-positive rate, and adapting the system for large-scale deployment in CUPID. Additional work still needs to be performed before the final CUPID muon veto design can be realized, including studies of prototype performance and longevity under the environmental conditions at the Laboratori Nazionali del Gran Sasso and consideration of any systematic effects that may arise from the muon distribution at Gran Sasso. In future work, we plan to address these uncertainties and the overall impact of all studies on the final CUPID muon veto design.
@@ Acknowledgments
We thank the technical staff of Wright Laboratory, particularly Tom Barker for his assistance in designing PCBs. We acknowledge the contributions of Penny Slocum, James Nikkel, Thomas Barker, Caitlin Gainey, Gabriel Hoshino, Jacky Hua, Neal Ma, Din-Ammar Tolj, Iffat Zarif, and Andrew Zheng to this research. The authors are appreciative of the members of the CUORE and CUPID collaborations for their support and expertise in this work. This paper includes data from a panel produced for a muon veto at the Low Background Counting Facility at the Lawrence Berkeley National Laboratory. We thank Rick Norman for providing this panel and feedback on this manuscript. We also thank Vincenzo Caracciolo for useful discussions and comments on the manuscript. Eljen Technologies, particularly Chris Maxwell, provided samples and extensive knowledge for this project, for which we are grateful. Kuraray and Saint-Gobain (now named Luxium Solutions) provided fiber samples and expertise. We would also like to thank Hamamatsu Photonics and CAEN Technologies for their technical support, particularly Yuri Venturini and Marco Locatelli. This work is supported by the DOE Office of Science, Office of Nuclear Physics, Contract No. DE-SC0012654. R.N.G. acknowledges grant NSF-PHY-1913374. This work made use of pandas [49], Matplotlib [50], Numpy [51], SciPy [52], and Jupyter [53].